$SPAD/src/input f02bjf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
n:=4
ia:=4
ib:=4
eps1:SF:=1.0e-4
matv:=true
iv:=4
a:Matrix SF:=
 [[3.9 ,12.5 ,-34.5 ,-0.5 ],_
  [4.3 ,21.5 ,-47.5 ,7.5 ],_
  [4.3 ,21.5 ,-43.5 ,3.5 ],_
  [4.4 ,26.0 ,-46.0 ,6.0 ]]
b:Matrix SF:=
 [[1 ,2 ,-3 ,1 ],_
  [1 ,3 ,-5 ,4 ],_
  [1 ,3 ,-4 ,3 ],_
  [1 ,3 ,-4 ,4 ]]
result:=f02bjf(n,ia,ib,eps1,matv,iv,a,b,-1)

———-

References^{[1] nothing}

4